OpenStudy (anonymous):

Ill fan and medal! What is the perimeter of a triangle with the given side lengths? s1 = x - 3 ft s2 = x + 2 ft s3 = x + 4 ft P=____ x + ____ ft. I don't understand how to figure this out. Plz help!

OpenStudy (anonymous):

@4Jesus @girlover @HaileyD @domofan @ClutzyMe21

OpenStudy (anonymous):

I have no clue as to how to figure this out....I'm super sorry! @summerforever21 May be able to help, idk

OpenStudy (anonymous):

@kpopaigoo @jabez177 @mrbaconqwerty @iRespect @angelic._.

OpenStudy (anonymous):

@nermaljean99 @holly01 @DUDE..LOOK..ITS..SGTC

OpenStudy (anonymous):

@micahm @brei @party_girl

OpenStudy (anonymous):

The perimeter is the sum of all the sides, correct? So just add the sides together \[P = ( x - 3 ) + ( x + 2 ) + ( x + 4 )\] The solution is the perimeter, which is the sum of all sides

OpenStudy (anonymous):

so would it be P= 3x+0? Im not good with math, and i get confused easily @holly01

OpenStudy (anonymous):

? @holly01

OpenStudy (anonymous):

s2 = x + 2 ft

OpenStudy (anonymous):

MEDAL!! PLZZ

OpenStudy (anonymous):

I need help with the answer... @_lilb_to_smooth_4_U_

OpenStudy (anonymous):

ok whats your question?

OpenStudy (anonymous):

look at the top. I have to figure out P=___x + ___ft

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

Does anyone know? i need help ASAP!!!

OpenStudy (anonymous):

9x is your answer http://widget.openstudy.com/study#/updates/54710687e4b001ceb26b1f99

OpenStudy (anonymous):

The correct answer was P= 3 + 3

OpenStudy (anonymous):

MEDAL?

OpenStudy (anonymous):

@girlsquirrel I'm so sorry i didn't reply to your question! If you have another question you would like me to answer i'll be glad to help. I deeply apologize about that

OpenStudy (anonymous):

that's ok! don't worry about it :) @holly01